OpenStudy (anonymous):

Simplify the expression of square root of 45?

OpenStudy (kropot72):

First step is to factorise 45. Can you find the factors of 45?

OpenStudy (anonymous):

1, 3, 5, 9, 15, and 45.

OpenStudy (anonymous):

The answer is 3.5 I think.

OpenStudy (kropot72):

\[\sqrt{45}=\sqrt{9 \times 5}\] One of these factors is a perfect square. Which one is it?

OpenStudy (kropot72):

What pairs of numbers multiplied give 9?